# JWST captures a sudden stellar outburst and inner disk wall destruction

### v1

@@ Abstract
We present JWST/MIRI observations of T Cha, a highly variable ($\Delta V\sim$3-5 mag) accreting Sun-like star surrounded by a disk with a large ($\sim 15$ au) dust gap. We find that the JWST mid-infrared spectrum is signiticantly different from the Spitzer spectrum obtained 17 years before, where the emission at short wavelengths ($5-10\mu m$) has decreased by $\sim 2/3$ while at longer wavelengths ($15-25\mu m$) it increased by up to a factor of $\sim 3$. This ‘seesaw’ behavior is contemporary with a fairly constant higher optical emission captured by the All Sky Automated Survey. By analyzing and modelling both SEDs, we propose that JWST caught the star during an outburst that destructed the asymmetric inner disk wall responsible for the high optical variability and lower $15-25$ micron emission during the Spitzer time. The dust mass lost during this outburst is estimated to be comparable ($\sim 1/5$) to the upper limit of the total micron-sized dust mass in the inner disk of T Cha now. Monitoring this system during possible future outbursts and more observations of its quiescent state will reveal if the inner disk can be replenished or will continue to be depleted and vanish.
facilities: JWST(MIRI MRS), Spitzer(IRS)software: astropy (Astropy Collaboration et al., 2013, 2018), JWST (Bushouse et al., 2023), scipy (Virtanen et al., 2020)
@@ 1 Introduction
Gas-rich dust disks around young stars are the birthplaces of planets, yet many questions on how planets form remain unanswered, highlighting the critical need for further study of their birth environments. These planet-forming disks are distinguished from Main Sequence stars by the infrared excess on top of the stellar emission. Among them, those having low near-infrared (NIR) excesses, but pronounced mid- and far-infrared (IR) excesses are particularly interesting. This type of spectral energy distribution (SED) indicates a discernible diminishment in dust mass in the inner disk and such disks are called transitional disks, as they are thought to be the transition stage from optically thick full disks to dissipated Class III (see Espaillat et al., 2014, for a review). While star-driven photoevaporation (e.g., Alexander et al., 2014) might explain the inside-out clearing of a subset of these objects, it is clear that those transition disks with large mass accretion rates and/or large dust cavities require a different process (see Fig. 6 in Ercolano & Pascucci, 2017). In a few cases planet formation is known to be the culprit of the dust clearing (e.g., Keppler et al., 2018; Christiaens et al., 2024)
@@
The central stars of the disks (also called T Tauri Stars) are young, accreting and variable at optical wavelengths. The Spitzer Space Telescope (hereafter Spitzer) revealed variability also at mid-IR wavelengths from their dust disks. An interesting subset of transition disks shows “seesaw” variability at mid-IR band, where the emission at shorter wavelengths varies inversely with the emission at longer wavelengths (e.g., Muzerolle et al., 2009; Espaillat et al., 2011; Flaherty et al., 2012; Kóspál et al., 2012). It was proposed that changes of the height of the optically thick inner edge or wall located close to the dust destruction radius (where the dust is hot and emits primarily at NIR wavelengths) could lead to the changes in shadowing of the outer disk (which emits at longer wavelengths), thus resulting in the seesaw effect (see Fig. 7 of Espaillat et al., 2014, for an illustration). The inner dust disk wall height variation can be caused by many factors including the dynamical interaction between the magnetic field and the disk (e.g., Bouvier et al., 1999; Lai & Zhang, 2008), perturbations in disks by planets, or turbulence (e.g., Espaillat et al., 2014). Besides inner disk wall height change, a warp in the disk can also result in seesaw changes in the SED (e.g., Nixon & Pringle, 2010).
@@
The launch of the James Webb Space Telescope (JWST) in December 2021 enables the analysis of the variabilities at mid-IR wavelengths again after Spitzer, including more long-term time ($\sim$ 15 years) variability of disks by combining both observations from both telescopes. In this paper, we present the JWST MIRI/MRS spectrum of a transition disk system around T Chamaeleontis (T Cha), a G8V T Tauri star (Alcala et al., 1993) located at a distance of $\sim$ 102.7 pc (Gaia Collaboration et al., 2021) in the $\epsilon$-Cha star forming region (Murphy et al., 2013), and compare it with a previously observed Spitzer IRS spectrum and other optical to millimeter photometric data. We show that the SED of T Cha has significantly changed compared to the Spitzer spectrum taken $\sim$17 years earlier. By modelling this change, we found that T Cha is likely to have lost its inner dust wall in 2021 and rebuilt it in 2023. We will show the observations of T Cha in section 2. The model to fit the observations is presented in section 3, and we will discuss the physical meaning and the implication of our model in section 4. We summarize our findings in section 5.
@@ 2 T Cha and its variable spectrum
T Cha has been long known to be surrounded by a circumstellar disk (e.g., Alcala et al., 1993, from the infrared excess of its SED). Brown et al. (2007) observed T Cha with the Spitzer and inferred a large dust gap ($\sim$15 au) from modeling its SED. The gap was confirmed by follow-up NIR interferometry (Olofsson et al., 2011, 2013) and high-resolution millimeter imagery (Hendler et al., 2018). The latter observations reveal the existence of an unresolved inner disk with radius $<$1 au, and an outer disk beyond $\sim$20 au with an inclination of $\sim$70. T Cha is also known to be accreting disk gas with an accretion rate of $\dot{M}\sim 1\times 10^{-9}$ to $3\times 10^{-8}M_{\odot}\,{\rm yr}^{-1}$ (Schisano et al., 2009; Cahill et al., 2019), suggesting that there is no significant gap in the gas (otherwise the mass accretion rate will be much lower), which is further supported by the ALMA CO gas emission map (Wölfer et al., 2023). Other stellar parameters of T Cha that are relevant for this study are summarized in Table 2.
Table 1: Properties of T Cha
@@
$\dot{M}$($M_{\odot}$/yr) 4$\times$10-9 4
T(K) 5600 5
References. — 1. Gaia Collaboration et al. (2021); 2. Alcala et al. (1993); 3. Olofsson et al. (2011); 4. Schisano et al. (2009) 5. Brown et al. (2007)
@@ 2.1 Variability of T Cha
Like many other young stellar objects, T Cha shows short-term photometric variability at optical and NIR wavelengths (e.g., Covino et al., 1992; Alcala et al., 1993; Walter et al., 2018). According to previous studies, the optical variability of T Cha amounts to a change in visual magnitude of $\sim 3$ (V magnitude from 10 to 13, e.g., Alcala et al., 1993; Walter et al., 2018). In this paper, we also incorporate optical monitoring from the All Sky Automated Survey (including ASAS3 during 2000-2009 and ASAS-SN from 2014), and find that at times T Cha can be as dim as $V\sim 15$ magnitude. This means that the total change in visual magnitude can be up to $\sim 5$ (see Fig 1). Such optical variability is much larger than the average T Tauri variability (e.g., Osterloh et al., 1996; Rigon et al., 2017). When optical emission varies, the emission from all bands (from B to K band) increase or decrease together, with redder colors when the emission fades (e.g., Walter et al., 2018). During the relatively less variable stage (when $\Delta V\sim 3$ mag, e.g., the green shaded region in Fig. 1), around half of the variability ($\Delta V\sim 1.5$ mag) shows a periodic behavior in $\sim$ 3.3 days (e.g., Mauder & Sosna, 1975; Walter et al., 2018), while there is also significant random scatter superimposed on the periodic variability (see Figure 5 of Walter et al., 2018).
Optical and NIR variability of young stellar objects can arise from many photospheric processes including outbursts, magnetic cool spots and unsteady accretion hot spots (e.g., Herbst et al., 1994; Fernandez & Eiroa, 1996; Fischer et al., 2023). Outbursts are typically one-time events and are not likely to be the origin of the short-term (a few hours to $\sim 3$ days) variability of T Cha (e.g., Lorenzetti et al., 2012; Findeisen et al., 2013; Fischer et al., 2023). Hot spots are typically caused by the accretion onto the star. They can lead to brighter emissions above the stellar photospheric model and are generally irregular on timescales as short as hours (e.g., Herbst et al., 1994), which can explain those photometric points that are randomly scattered. However, periodic variations in the stellar photosphere like cool spots are typically weak (largest $\Delta V\sim 0.8$ mag), and are unable to explain the periodic optical variability seen in most epochs of T Cha ($\Delta V\gtrsim 1.5$ mag). In addition, the $\sim$ 3.3 days variability is consistent with the period of a Keplerian orbit at 0.04--0.05 au. This can be considered as the corotation radius and is close to magnetic truncation radius ($\sim 5-10~{}R_{*}$, e.g., Gullbring et al., 1998; Bouvier et al., 2007), which is the inner edge of the gas disk. Thus, previous studies suggest additional extinction caused by the circumstellar disk at this radius, such as a inner disk wall, is required (e.g., Herbst et al., 1994; Schisano et al., 2009; Walter et al., 2018) and is also plausible due to the high disk inclination (e.g., Pohl et al., 2017; Hendler et al., 2018).
Figure 1: Optical and mid-IR light curves of T Cha. The lines and shaded area indicates the date when some of the photometric points we analyzed in this paper were taken. Blue shaded areas indicates the period when T Cha has less variable but brighter optical emission. The ‘LCOGT and Andicam’ is from Walter et al. (2018).
@@
Recent WISE measurements, however, reveal a drop in flux of T Cha at 3 and 5 $\mu$m in 2015 and 2021 (by a factor of $\sim 40\%$ and $\sim 70\%$ in flux, respectively). Both drops are accompanied by a brighter and less variable optical emission [^1] Similar optical behavior was also seen in 1989 (Covino et al., 1992; Alcala et al., 1993) and 2006 (blue shaded area in Fig. 1), but no contemporary IR observation was taken. seen in ASAS surveys (Fig. 1). The larger change in 2021 also persists for longer, with both the decrease in WISE bands and the constant brighter optical emission lasting for $\sim 2$ years. Observations from The American Association of Variable Star Observers (AAVSO) shown in Fig 2 identified the same optical and NIR behavior in different bands. Comparing the emission during the 2021 optically brighter period with historical photometry (e.g., Covino et al., 1992; Alcala et al., 1993) shows that T Cha gets bluer when it gets brighter, implying less extinction and possibly more accretion during these periods. The change in both optical and IR bands indicates a change in both the inner disk (e.g., less extinction) and the stellar photospheric processes (e.g., accretional hot spots or outbursts).
Figure 2: Left Panel: AAVSO photometric data (colored) plotted with the ASAS-SN data (gray). The JWST observation time and blue shaded area are plotted the same as Fig 1. The emissions of T Cha in all optical and NIR bands are less variable and brighter during the same period between 2021 and 2023. Right Panel: Color magnitude diagram from AAVSO data in the blue shaded period (circles) compared with the photometry during 1991-1992 (Alcala et al., 1993) (crosses) and 1989 (Covino et al., 1992) (dimonds). When T Cha gets brighter, it gets bluer.
@@ 2.2 JWST MIRI/MRS reveals mid-IR changes for T Cha
Figure 3: Overview of the SED of T Cha. The grey dased line is 5600 K stellar photosphere emission. The red and blue solid lines are Spitzer/IRS SED and JWST MIRI/MRS spectra, respectively. The two strong lines in the JWST spectrum are [ArII] and [NeII], and the broad features are PAH features, which are discussed in Bajaj et al. (2024). The photometric points in optical bands are from Alcala et al. (1993); Walter et al. (2018) and AAVSO, and the IR photometric points are from various papers (see Table 5). We highlight some points in blue to indicate they are taken contemporary with the JWST observation (see the blue shaded region next to the JWST line in Fig 1). The near and mid-IR points connected with dotted lines indicate one contemporary observation at JHKLM bands (Alcala et al., 1993).
@@ 3 Models and Analyses
In this section, we discuss disk models that can reproduce the overall SED of T Cha as well as the change between the Spitzer/IRS and the JWST MIRI-MRS spectra. Specifically, these models= need to account for: 1. The mid-IR seesaw behavior, which indicates less emission from the hot inner disk and more emission from the cool outer disk, from the Spitzer to the JWST spectrum. 2. The periodic part of the optical variability which is too strong to be only due to stellar variability, and is significantly reduced in 2021.
@@ 3.1 Model setup
We use the DiskMINT wrapper (Deng et al., 2023) to model the SED of T Cha with the radiative transfer code RADMC-3D (Dullemond et al., 2012). In our model, we adopt the stellar radius and mass $R_{\star}=1.3R_{\odot}$, $M_{*}=1.5M_{\odot}$ from Alcala et al. (1993); Olofsson et al. (2011). We adopt the stellar photospheric spectrum from the 5600 K [^2] We note that the temperature of young stellar objects can be influenced by many factors. We tried with a temperature range of 5100-5900K, and it doesn’t affect the main points discussed in this paper. (e.g., Brown et al., 2007) BT-Settl (AGSS2009) (e.g. Allard et al., 2003, 2011) with the assumption of solar metallicity. To run RADMC-3D we smooth the input stellar photospheric model with scipy.interpolate.interp1d and scipy.signal.savgol_filter to remove absorption lines, making sure that the bolometric luminosity is conserved. We use the dsharp_opac package from Birnstiel et al. (2018) to generate the input opacity file of the dust incorporating constants from Draine (2003) for graphite and Weingartner & Draine (2001) for astronomical silicate. The best-fit model is found by minimizing the $\chi^{2}$ between the observed mid-IR SED and that derived from our model.
@@ 3.1.1 Disk structure
Before the JWST observation, several studies have focused on modelling the SED of T Cha and have provided good constraints on the outer disk (e.g., Brown et al., 2007; Olofsson et al., 2013; Huélamo et al., 2015). However, previous models have not taken the optical variability into account and assume an azimuthally symmetric inner disk. Inspired by the hypothesis that the inner disk blocking the stellar photosphere is the cause of short-term optical variability (e.g., Covino et al., 1992; Herbst et al., 1994), we adopted an asymmetric inner disk model to fit the Spitzer spectrum along with the optical variability. We consider that all the mass in the inner disk is concentrated in $\delta\phi$ by setting the mass in other regions to zero (see Fig 4 for an illustration), and ignore the effect of dynamics caused by heating and cooling. In this way, when the optically thick part rotates between the star and our line of sight, we will see a highly extincted stellar photosphere, and when it rotates away there will be little/no extinction. On the other hand, for the less variable bright optical points between 2021 and 2023, extinction from the inner disk wall is not necessary. For simplicity, we use an azimuthally symmetric inner disk to model the SED at the time of the JWST observations. We note that the $\delta\phi$ also represents an azimuthal concentration of inner disk wall and can significantly influence the SED, thus it’s highly degenerate with other parameters in SED fitting (see appendix A for more details). Thus we sample $\delta\phi$ coarsely (from $0^{\circ}$ to $360^{\circ}$ with a step of $60^{\circ}$) to simplify our modelling while showing the trend of how it influences both the optical and IR emissions.
@@
Since almost all of the disk emission comes from the inner wall and the outer disk (e.g., Olofsson et al., 2013, and in our model discussed in the following sections), we add the hidden mass component only for estimating the upper limit of the micron-sized dust mass in the inner disk, thus its azimuthal distribution also follows the inner disk wall. The amount of mass in this component is an upper limit because if more mass is added it will contribute to the mid-IR emission resulting in a peak at $\sim 9\mu m$, which is not consistent with our observation. The ALMA image constrains the inner disk of T Cha to be smaller than 1 au (Hendler et al., 2018), which we set as the outer radius of the hidden mass. Other parameters of the hidden mass are set or sampled the same as the inner disk wall, with the scale height limited to be lower than the wall. How the fixed parameters in Table 2 are chosen is discussed in detail in the following section 3.1.2.
Figure 4: Sketch of the disk structure of our model. Sizes are not to scale. We fit three components: the inner disk dust wall, the hidden mass behind the wall and the outer disk together. The inner disk components can be azimuthally asymmetric.

### v3

@@ Abstract
We analyze JWST/MIRI observations of T Cha, a highly variable ($\Delta V\sim$3-5 mag) accreting Sun-like star surrounded by a disk with a large ($\sim 15$ au) dust gap. We find that the JWST mid-infrared spectrum is significantly different from the Spitzer spectrum obtained 17 years before -- the emission at short wavelengths ($5-10\mu m$) has decreased by $\sim 2/3$ while that at longer wavelengths ($15-25\mu m$) has increased by up to a factor of $\sim 3$. The JWST spectrum is contemporary with a fairly constant higher optical emission captured by the All Sky Automated Survey. By analyzing and modelling both SEDs, we propose that JWST caught the star during an outburst that partly destroyed and significantly reduced the height of the asymmetric inner disk wall responsible for the high optical variability and lower $15-25$ µm emission during the Spitzer time. The dust mass lost during this outburst is estimated to be comparable ($\sim 1/5$) to the upper limit of the total micron-sized dust mass in the inner disk of T Cha now. Monitoring this system during possible future outbursts and more observations of its quiescent state will reveal if the inner disk can be replenished or will continue to be depleted and vanish.
facilities: JWST(MIRI MRS), Spitzer(IRS)software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), JWST (Bushouse et al., 2023), scipy (Virtanen et al., 2020)
@@ 1 Introduction
Gas-rich dust disks around young stars are the birthplaces of planets, yet many questions on how planets form remain unanswered, highlighting the critical need for further studies of their birth environments. These planet-forming disks were first identified via infrared excess emission on top of the stellar photosphere. Among them, those having low near-infrared (NIR) excess, but pronounced mid- and far-infrared (IR) excess are particularly interesting as this type of spectral energy distribution (SED) indicates a discernible diminishment in dust mass in the inner disk. Such disks are called transitional disks and thought to be the transition stage from optically thick full disks to dissipated Class III sources (see Espaillat et al., 2014, for a review). While star-driven photoevaporation (e.g., Alexander et al., 2014) might explain the inside-out clearing of a subset of these objects, it is clear that those transition disks with large mass accretion rates and/or large dust cavities require a different process (see Fig. 6 in Ercolano & Pascucci, 2017). In a few cases planet formation is known to be the culprit of the dust clearing (e.g., Keppler et al., 2018; Christiaens et al., 2024)
@@
The central stars of the disks (also called T Tauri Stars) are young, accreting and variable at optical wavelengths. The Spitzer Space Telescope (hereafter Spitzer) revealed variability also at mid-IR wavelengths from their dust disks. An interesting subset of transition disks shows “seesaw” variability at mid-IR band, where the emission at shorter wavelengths varies inversely with the emission at longer wavelengths (e.g., Muzerolle et al., 2009; Espaillat et al., 2011; Flaherty et al., 2012; Kóspál et al., 2012; Espaillat et al., 2024). It was proposed that changes of the height of the optically thick inner edge or wall located close to the dust destruction radius (where the dust is hot and emits primarily at NIR wavelengths) could lead to the changes in shadowing of the outer disk (which emits at longer wavelengths), thus resulting in the seesaw effect (see Fig. 7 of Espaillat et al., 2014, for an illustration). The inner dust disk wall height variation can be caused by many factors including the dynamical interaction between the stellar magnetic field and the disk (e.g., Bouvier et al., 1999; Lai & Zhang, 2008), perturbations in disks by planets, or turbulence (e.g., Espaillat et al., 2014). Besides changes in the inner disk wall height, a warp in the disk can also result in seesaw changes in the SED (e.g., Nixon & Pringle, 2010).
@@
The launch of the James Webb Space Telescope (JWST) in December 2021 opens new opportunities for investigating variability at mid-IR wavelengths again after Spitzer, including more long-term ($\sim$ 15 years) variability of disks by combining observations from both telescopes. In this paper, we present an analysis of the JWST MIRI/MRS spectrum of a transition disk system around T Chamaeleontis (hereafter T Cha, see Bajaj et al., 2024, for details about the observations). T Cha is a G8V T Tauri star (Alcala et al., 1993) located at a distance of $\sim$ 102.7 pc (Gaia Collaboration et al., 2021) in the $\epsilon$-Cha star forming region (Murphy et al., 2013). We compare the 2022 MIRI spectrum with a previous Spitzer IRS spectrum obtained in 2005 and other optical to millimeter photometric data. We show that the SED of T Cha has significantly changed compared to the Spitzer epoch, $\sim$17 years earlier. By modelling this change along with ASAS-SN monitoring, we found that T Cha is likely to have partly lost its inner dust wall in 2021 and regained it in 2023. The SED of T Cha is discussed in section 2. The model to fit the observations is presented in section 3, while the physical meaning and the implications of our model are outlined in section 4. We summarize our findings in section 5.
@@ 2 T Cha and its variable spectrum
T Cha has been long known to be surrounded by a circumstellar disk (e.g., Alcala et al., 1993). Brown et al. (2007) observed T Cha with Spitzer and inferred a large dust gap ($\sim$15 au) from modeling its SED. The gap was confirmed by follow-up NIR interferometry (Olofsson et al., 2011, 2013) and high-resolution millimeter imagery (Hendler et al., 2018). The latter observations reveal the existence of an unresolved inner disk with radius $<$1 au, and an outer disk beyond $\sim$20 au with an inclination of $\sim$70. T Cha is also known to be accreting disk gas with an accretion rate ($\dot{M}$) varying from $\sim 1\times 10^{-9}$ to $3\times 10^{-8}M_{\odot}\,{\rm yr}^{-1}$ (Schisano et al., 2009; Cahill et al., 2019). Continued accretion and 12CO gas emission maps obtained with ALMA (Wölfer et al., 2023) suggest that there is no significant gap in the gas. Stellar parameters of T Cha that are relevant for this study are summarized in Table 2.
Table 1: T Cha stellar properties relevant to our study
@@
[…]
T(K) 5600 6
References. — 1. Gaia Collaboration et al. (2021); 2. Alcala et al. (1993); 3. Olofsson et al. (2011); 4. Schisano et al. (2009) 5.  Cahill et al. (2019) 6. Brown et al. (2007)
@@ 2.1 Variability of T Cha
Like many other young stellar objects, T Cha shows short-term photometric variability at optical and NIR wavelengths (e.g., Covino et al., 1992; Alcala et al., 1993; Walter et al., 2018). According to previous studies, the optical variability of T Cha amounts to a change in visual magnitude of $\sim 3$ (V magnitude from 10 to 13, e.g., Alcala et al., 1993; Walter et al., 2018). In this paper, we also incorporate optical monitoring from the All Sky Automated Survey (including ASAS3 during 2000-2009 and ASAS-SN from 2014), and find that at times T Cha can be as dim as $V\sim 15$ mag. This means that the total change in visual magnitude can be up to $\sim 5$ (see Fig 1). Such optical variability is much larger than the average T Tauri variability (e.g., Osterloh et al., 1996; Rigon et al., 2017). About half of the optical variability shows a periodic behavior in $\sim$ 3.3 days (e.g., Mauder & Sosna, 1975; Walter et al., 2018), while there is also significant random scatter superimposed on the periodic variability. The emission from all bands (from B to K band) increases or decreases together, with redder colors when the emission fades (e.g., Walter et al., 2018, and Figure 2).
@@
Figure 1: Optical and mid-IR light curves of T Cha. The lines and shaded areas indicate the date when some of the photometric points we analyzed in this paper were taken. Blue shaded areas indicate the period when T Cha has less variable but brighter optical emission.
[…]
Recent WISE measurements, however, reveal a drop in flux of T Cha at 3 and 5 $\mu$m in 2015 and 2021 (by a factor of $\sim 40\%$ and $\sim 70\%$ in flux, respectively). Both drops are accompanied by a brighter and less variable optical emission [^1] Similar optical behavior was also seen in 1989 (Covino et al., 1992; Alcala et al., 1993) and 2006 (blue shaded area in Fig. 1), but no contemporary IR observation was taken. seen in ASAS surveys (Fig. 1). The larger change in 2021 also persists for longer, with both the decrease in WISE bands and the constant brighter optical emission lasting for $\sim 2$ years. Observations from The American Association of Variable Star Observers (AAVSO) shown in Fig 2 identified the same optical and NIR behavior in different bands and comparison with historical photometry (e.g., Covino et al., 1992; Alcala et al., 1993) positions T Cha during this time on the brightest optical points and smallest color indices. The change in both optical and IR bands indicates a possible change in both the inner disk (e.g., less extinction) and the stellar photosphere (e.g., more accretion or an outburst).
Figure 2: Left Panel: AAVSO photometric data (colored) plotted with the ASAS-SN data (gray). The JWST observation time and blue shaded area are plotted the same as Fig 1. The emissions of T Cha in all optical and NIR bands are less variable and brighter during the same period between 2021 and 2023. Right Panel: Color magnitude diagram from AAVSO data in the blue shaded period (circles) compared with the photometry during 1991 (Alcala et al., 1993) (crosses) and 1989-1990 (Covino et al., 1992) (dimonds). When T Cha gets brighter, it gets bluer.
@@ 2.2 JWST MIRI/MRS reveals mid-IR changes for T Cha
Figure 3: Overview of the SED of T Cha. The grey dased line is 5600 K stellar photosphere emission scaled to the distance to T Cha. The red and blue solid lines are Spitzer/IRS SED and JWST MIRI/MRS spectra, respectively. The two strong lines in the JWST spectrum are [ArII] and [NeII], and the broad features are PAH features. The photometric points in optical bands are from Alcala et al. (1993); Walter et al. (2018) and AAVSO, and the IR photometric points are from various papers (see Table 5). We highlight some points in blue to indicate they are taken contemporary with the JWST observation (see the blue shaded region next to the JWST line in Fig 1). The near and mid-IR points connected with dotted lines indicate one contemporary observation at JHKLM bands (Alcala et al., 1993).
@@ 3 Models and Analyses
In this section, we discuss disk models that can reproduce the overall SED of T Cha as well as the change between the Spitzer/IRS and the JWST MIRI-MRS spectra. Specifically, these models need to account for: 1. The mid-IR seesaw behavior, which indicates less emission from the hot inner disk and more emission from the cool outer disk, from the Spitzer to the JWST spectrum. 2. The periodic part of the optical variability which is too strong to be only due to stellar variability, and is significantly reduced in 2021.
@@ 3.1 Model setup
We use the DiskMINT wrapper (Deng et al., 2023) to model the SED of T Cha with the radiative transfer code RADMC-3D (Dullemond et al., 2012). In our model, we adopt the stellar radius and mass $R_{\star}=1.3R_{\odot}$, $M_{*}=1.5M_{\odot}$ from Alcala et al. (1993); Olofsson et al. (2011). We adopt the stellar photospheric spectrum from the 5600 K [^2] We note that the temperature of young stellar objects can be influenced by many factors. We tried with a temperature range of 5100-5900K, and it doesn’t affect the main points discussed in this paper. (e.g., Brown et al., 2007) BT-Settl (AGSS2009) (e.g. Allard et al., 2003, 2011) with the assumption of solar metallicity. The accretion shock term is hard to determine for T Cha due to the lack of contemporary FUV data, and has little impact on the mid-IR modelling and our main results. Thus, we ignore this term for simplicity. To run RADMC-3D we smooth the input stellar photospheric model with scipy.interpolate.interp1d and scipy.signal.savgol_filter to remove absorption lines, making sure that the bolometric luminosity is conserved. We use the dsharp_opac package from Birnstiel et al. (2018) to generate the input opacity file of the dust incorporating constants from Draine (2003) for graphite and Weingartner & Draine (2001) for astronomical silicate. With the input stellar photosphere and the disk model with opacity file, the code performs radiative transfer calculations to generate the temperature structure and the SED. The best-fit model is found by minimizing the $\chi^{2}$ between the observed mid-IR SED and that derived from our model.
@@ 3.1.1 Disk structure
Before the JWST observation, several studies have focused on modelling the SED of T Cha and have provided good constraints on the outer disk (e.g., Brown et al., 2007; Olofsson et al., 2013; Huélamo et al., 2015). However, previous models have not taken the optical variability into account and assume an azimuthally symmetric inner disk. Inspired by the hypothesis that the inner disk blocking the stellar photosphere is the cause of short-term optical variability (e.g., Covino et al., 1992; Herbst et al., 1994), we adopted an asymmetric inner disk model to fit the Spitzer spectrum along with the optical variability. We consider that all the mass in the inner disk is concentrated in $\delta\phi$ by setting the mass in other regions to zero (see Fig 4 for an illustration), and ignore the effect of dynamics caused by heating and cooling. In this way, when the optically thick part rotates between the star and our line of sight, we will see a highly extincted stellar photosphere, and when it rotates away there will be little/no extinction. On the other hand, for the less variable ($\delta V<0.2$ mag in most cases) bright optical points between 2021 and 2023, extinction from the inner disk wall is not necessary. For simplicity, we use an azimuthally symmetric inner disk to model the SED at the time of the JWST observations. We note that the $\delta\phi$ also represents an azimuthal concentration of inner disk wall and can significantly influence the SED, thus it is highly degenerate with other parameters in SED fitting (see appendix A for more details). As such, we sample $\delta\phi$ coarsely (from $0^{\circ}$ to $360^{\circ}$ with a step of $60^{\circ}$) to simplify our modelling while showing the trend of how it influences both the optical and IR emissions.
@@
Since almost all of the disk emission comes from the inner wall and the outer disk (e.g., Olofsson et al., 2013, and in our model discussed in the following sections), we add the hidden mass component only for estimating the upper limit of the micron-sized dust mass in the inner disk, thus its azimuthal distribution also follows the inner disk wall. The amount of mass in this component is an upper limit because if more mass is added it will no longer be hidden and contribute to the mid-IR emission resulting in a silicate emission feature at $\sim 10$ micron, which is not consistent with the observations. The ALMA image constrains the inner disk of T Cha to be smaller than 1 au (Hendler et al., 2018), which we set as the outer radius of the hidden mass. Other parameters of the hidden mass are set or sampled the same as the inner disk wall, with the scale height limited to be lower than the wall. How the fixed parameters in Table 2 are chosen is discussed in detail in the following section 3.1.2.
Figure 4: Sketch of the dust disk structure of our model. Sizes are not to scale. We fit three components: the inner disk dust wall, the hidden mass behind the wall and the outer disk together. The inner disk components can be azimuthally asymmetric.